A chemist has 35% NaCl solution and 50% NaCl solution. How many liters of each should he mix to get 10L of NaCl solution with 40% base concentration?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

A chemist has 35% NaCl solution and 50% NaCl solution. How many liters of each should he mix to get 10L of NaCl solution with 40% base concentration?
---------------
Equations:
Quantity: x + y = 10L
Quality: 0.35x + 0.50y = 0.40*10
-------------------------------------
Multiply thru 1st by 25
Multiply thru 2nd by 100
-------------------------------
Quantity: 35x + 35y = 35*10
Quality: 35x + 50y = 40*10
---------------------------------
Subtract Quantity from Quality to get:
15y = 5*10
y = 3 1/3 L (amt of 50% NaCl needed in the mixture)
---
Solve for "x":
x = 10- 3 1/3 = 6 2/3 L (amt of 35% NaCl needed in the mixture)
===============
Cheers,
Stan H.

Answer by mananth(16946)

35% NaCl solution let this be x L
50% NaCl solution. this will be 10-x
mix to get 40% 10 L
..
0.35x+0.5(10-x)=0.4*10
0.35x+5-0.5x=4
-0.15x=-1
x=-1/0.15
= 6.66 L 35% solution
Balance 3.34 L will be 50% solution
CHECK
6.66*35+3.34*50 = 40*10